# Towards Sequence Modeling Alignment between Tokenizer and Autoregressive Model

### v2

@@ Abstract
Autoregressive image generation aims to predict the next token based on previous ones. However, this process is challenged by the bidirectional dependencies inherent in conventional image tokenizations, which creates a fundamental misalignment with the unidirectional nature of autoregressive models. To resolve this, we introduce AliTok, a novel Aligned Tokenizer that alters the dependency structure of the token sequence. AliTok employs a bidirectional encoder constrained by a causal decoder, a design that compels the encoder to produce a token sequence with both semantic richness and forward-dependency. Furthermore, by incorporating prefix tokens and employing a two-stage tokenizer training process to enhance reconstruction performance, AliTok achieves high fidelity and predictability simultaneously. Building upon AliTok, a standard decoder-only autoregressive model with just 177M parameters achieves a gFID of 1.44 and an IS of 319.5 on the ImageNet-256 benchmark. Scaling up to 662M parameters, our model reaches a gFID of 1.28, surpassing the state-of-the-art diffusion method while achieving a 10 faster sampling speed. The code and weights are available at AliTok.
@@ 1 Introduction
Autoregressive models, particularly GPT-style decoder-only transformers (Achiam et al., 2023; Touvron et al., 2023a; b), have achieved revolutionary success in natural language processing owing to their simple yet scalable next-token prediction paradigm. Inspired by this triumph, the research community is actively exploring the application of this powerful paradigm to image generation (Han et al., 2024; Team et al., 2025; Tang et al., 2025; Xiong et al., 2025; Tian et al., 2024; Wang et al., 2024) by sequentially predicting a stream of tokens, showcasing a promising path for multi-modal unification (Team, 2024; Wu et al., 2025; Chen et al., 2025; Xie et al., 2024; Hurst et al., 2024).
However, as demonstrated in previous literature (Pang et al., 2024; Li et al., 2024), the inherent bi-directional property of visual sequences makes it difficult for the raster-scan decoder-only autoregressive model (e.g., LlamaGen (Sun et al., 2024)) to achieve excellent performance, since the unidirectional transformer is hard to model bi-directional sequences. Consequently, recent works have shifted focus towards paradigms like masked autoregressive (Li et al., 2024; Chang et al., 2022; Yu et al., 2024c; Fan et al., 2024) and next-scale prediction (Tang et al., 2024; Tian et al., 2024; Han et al., 2024), which employ bidirectional attention in autoregressive models and demonstrate a superior choice for visual modeling. However, these approaches complicate visual generation and diverge from the traditional autoregressive paradigm, increasing the challenges for multi-modal unification. In contrast to existing works that modify the model to fit the data properties, we propose an alternative perspective: Can we instead instill a forward-dependency into the token sequence, to align with the powerful simplicity of decoder-only AR models?
@@ 2 Related work
Autoregressive visual generation. Inspired by GPT-style language models (Achiam et al., 2023; Touvron et al., 2023a; b), early visual autoregressive methods (Van Den Oord et al., 2017; Lee et al., 2022) predict image tokens along the raster-scan order. However, this strict unidirectional paradigm struggles to effectively model the bidirectional spatial dependencies inherent in visual sequence, thereby limiting its performance ceiling. To resolve this fundamental conflict, research has pivoted to re-engineering the models or generation paradigms for stronger bidirectional context awareness. One prominent approach is masked autoregressive modeling, including MaskGIT (Chang et al., 2022) and MAR (Li et al., 2024), which leverages bidirectional attention for multi-round parallel generation. Another category explores hierarchical or next-scale prediction, such as VAR (Tian et al., 2024), which proceeds autoregressively across scales yet bidirectionally within each scale. Furthermore, other efforts enhance global perception by maximizing all-directional generation probabilities (RAR, Yu et al. (2024b)) or adopting random generation orders (RandAR, Pang et al. (2024)). While effective, these approaches compromise the simplicity of the classical autoregressive paradigm by introducing complex generation mechanisms or training objectives. Unlike these strategies that adapt models to data, this paper focuses on reshaping the data itself. We propose a novel tokenizer that instills a causal dependency structure into the token sequence during the encoding stage, thereby enabling a standard decoder-only autoregressive model to unleash its full potential.
Image tokenizer, a type of Variational Autoencoder (VAE) (Kingma et al., 2013), plays an important role in visual generation domain (Wan et al., 2025; Wu et al., 2024). By mapping high-dimensional image pixels to a low-dimensional latent space, the image tokenizer significantly enhances the training efficiency of generative models. Among them, VQ-VAE (Van Den Oord et al., 2017) utilizes quantization techniques to discretize continuous latent features for autoregressive image generation. Subsequently, VQ-GAN (Esser et al., 2021) substantially improves visual fidelity by incorporating an adversarial loss. MagViT-v2 (Yu et al., 2024a) and FSQ (Mentzer et al., 2023) focus on algorithmic improvements to efficiently scale the codebook size. Besides, TiTok (Yu et al., 2024c) proposes a transformer-based tokenizer, encoding 2D image patches into a 1D sequence. FlexTok (Bachmann et al., 2025) explores multi-granularity semantic encoding, ranging from coarse to fine-grained. Despite numerous efforts, there has been limited exploration of how to design a tokenizer that simultaneously achieves high-quality reconstruction while being more conducive to subsequent autoregressive generation. Bridging this critical gap is the core goal of our work.
@@ 3.2 AliTok tokenizer
The reconstruction loss for the first stage, follows standard practices (Yu et al., 2024c), combining a mean squared error (MSE) loss , perceptual loss , quantization loss (Zheng & Vedaldi, 2023) , and adversarial loss :
@@
Retraining a bidirectional decoder. In the second stage, we freeze the encoder and codebook of the tokenizer and retrain a bidirectional decoder to improve detail consistency. As shown in Fig. 4(d), in addition to converting the attention mechanism in the decoder, we integrate 64 buffer tokens (Li et al., 2024) to enhance modeling capabilities by increasing computational load. Meanwhile, the loss previously imposed on prefix tokens is removed to allow the decoder to focus on reconstruction quality. This two-stage training strategy enables our AliTok tokenizer to not only produce generation-friendly encoded tokens, but also maintain good reconstruction performance.
@@ 3.3 Autoregressive model
Our decoder-only autoregressive architecture follows the standard design in LlamaGen (Sun et al., 2024), which uses RMSNorm (Zhang & Sennrich, 2019) for pre-normalization and applies 2D rotary positional embeddings (RoPE) (Su et al., 2024) at each layer. Building upon LlamaGen, we make only a few modifications. First, since our method has 16 extra prefix tokens and needs to model 272 tokens, 2D RoPE cannot be used directly. Therefore, we use 1D RoPE for the prefix tokens and 2D RoPE for the remaining 256 tokens. Secondly, we introduce the QK-Norm operation (Team, 2024) in the attention module to stabilize the large-scale model training, aligning with recent multi-modal autoregressive models (Xie et al., 2024; Ma et al., 2025; Team, 2024). We do not conduct extensive exploration of autoregressive models and only use the conventional architecture to verify the effectiveness of the proposed tokenizer.
@@ 4.1 Experimental Setup
Models and evaluations. The design of the proposed tokenizer is based on TA-TiTok (Kim et al., 2025), utilizing ViT-B (Yu et al., 2024c) for the encoder and ViT-L for the decoder. We set the vocabulary size to 4096 and adopt the online feature clustering method (Zheng & Vedaldi, 2023) to ensure that the codebook utilization is 100%. For simplicity, we call the autoregressive models trained using our AliTok tokenizer as AliTok-B/L/XL, comprising three different sizes, more details are provided in Appendix A.4. Following common practice, FID (including rFID and gFID) (Heusel et al., 2017), IS (Salimans et al., 2016), Precision, and Recall are adopted as metrics. We generate 50,000 images to test the gFID and adopt the evaluation code from RAR (Yu et al., 2024b). For rFID, we use the ImageNet-1K val set for evaluation, consistent with other methods (Sun et al., 2024; Yu et al., 2024c; Yao et al., 2025). During the test, KV-cache is employed to enhance sampling speed.
Dataset and training details. We train our image tokenizer from scratch on ImageNet-1K (Russakovsky et al., 2015). The tokenizer is trained for 600K steps in the first stage and 300K steps in the second stage on 32 A800-80G GPUs. For the autoregressive model training, we employ tencrop (Szegedy et al., 2015) for data augmentation and cache (Li et al., 2024) the encoded results of the tokenizer to reduce training time. We train the base model and large model for 800 epochs and the XL model for 400 epochs, with a batch size of 2048 and a learning rate of 4e-4. For the first 100 epochs, the learning rate is linearly warmed up, then decayed using a cosine decay schedule down to 1e-5, following RAR (Yu et al., 2024b). During training, class conditioning is randomly dropped with a probability of 0.1 to support the use of classifier-free guidance (cfg) (Ho & Salimans, 2022).
@@ 4.2 Main Results
Sampling efficiency comparison. Beyond its excellent generation quality, AliTok also demonstrates a significant advantage in sampling efficiency benefiting from the use of KV-cache. As shown in Table 2, with similar gFID scores, AliTok-L improves throughput by 33.7 and 2.0 compared to MAR-H (Li et al., 2024) and RAR-XXL (Yu et al., 2024b), respectively. Compared to the state-of-the-art diffusion model LightningDiT (Yao et al., 2025), our AliTok-XL requires less than 10% of the time needed by LightningDiT to produce an image, establishing a substantial improvement in image generation efficiency.
Generation visualization. We further show the generation results under different model sizes in Fig. 5. It can be observed that even the smallest model is capable of producing high-quality visual results and reasonably generating complex structures, such as the coral reef and altar. As the model size increases, the generated images exhibit enhanced detail and more intricate structures, illustrated by the realistic texture of the feathers of the parrot, the fine-grained structure of the space shuttle and the altar, demonstrating a superior level of realism in image quality. More results generated by the different model sizes are provided in the Appendix (Fig. 9, 12, 13, 14, and 15).
@@
Reconstruction performance. We report the reconstruction performance of various tokenizers (Yu et al., 2024c; Zheng & Vedaldi, 2023; Xiong et al., 2025) in Table 4. By introducing a powerful bidirectional decoder in the second stage, AliTok significantly improves the rFID to 0.86. This result remains competitive with GigaTok-B-L (rFID: 0.81), particularly given that the latter relies on a substantially larger vocabulary size (16384 4096). Experiments demonstrate that our proposed AliTok successfully achieves an ideal balance between the goals of generation-friendly and high-fidelity through its ingenious decoupling design.
@@ A.1 Limitation and future work
Benefiting from the AliTok tokenizer, our 662M AR model reaches a significantly high training accuracy of 12.2%. In this context, the reconstruction performance of the discrete tokenizer becomes a critical bottleneck for the generative performance. To test the limiting value of gFID that can be achieved under this tokenizer, we randomly select 50 real images per class from the training set and calculate the gFID after reconstruction (differently from rFID evaluation), obtaining a FID of 1.15. This implies that even if the accuracy in the generation phase is close to 100% (which is not possible), the generation performance still cannot reach the upper limit of 1.15. This experimental result explains the phenomenon observed in Fig. 6, where the training loss and error rate of AliTok-XL continue to decrease steadily, yet the gFID metric struggles to achieve further improvement. Therefore, this paper does not explore larger generative models, and leaves the exploration of training a larger codebook, e.g., 16384 in LlamaGen (Sun et al., 2024), for an AliTok-like generation-friendly tokenizer to future work.

### v3

@@ Abstract
Autoregressive image generation aims to predict the next token based on previous ones. However, this process is challenged by the bidirectional dependencies inherent in conventional image tokenizations, which creates a fundamental misalignment with the unidirectional nature of autoregressive models. To resolve this, we introduce AliTok, a novel Aligned Tokenizer that alters the dependency structure of the token sequence. AliTok employs a bidirectional encoder constrained by a causal decoder, a design that compels the encoder to produce a token sequence with both semantic richness and forward-dependency. Furthermore, by incorporating prefix tokens and employing a two-stage tokenizer training process to enhance reconstruction performance, AliTok achieves high fidelity and predictability simultaneously. Building upon AliTok, a standard decoder-only autoregressive model with just 177M parameters achieves a gFID of 1.44 and an IS of 319.5 on ImageNet-256. Scaling to 662M, our model reaches a gFID of 1.28, surpassing the SOTA diffusion method with 10 faster sampling. On ImageNet-512, our 318M model also achieves a SOTA gFID of 1.39. Code and weights at AliTok.
@@ 1 Introduction
Autoregressive models, particularly GPT-style decoder-only transformers (Achiam et al., 2023; Touvron et al., 2023a; b), have achieved revolutionary success in natural language processing owing to their simple yet scalable next-token prediction paradigm. Inspired by this triumph, the research community is actively exploring the application of this powerful paradigm to image generation (Team et al., 2025; Han et al., 2024; Wei et al., 2025) by sequentially predicting a stream of tokens, showcasing a promising path for multi-modal unification (Team, 2024; Wu et al., 2025; Chen et al., 2025).
However, as demonstrated in previous literature (Pang et al., 2024; Li et al., 2024b), the inherent bi-directional property of visual sequences makes it difficult for the raster-scan decoder-only autoregressive model (e.g., LlamaGen (Sun et al., 2024)) to achieve excellent performance, since the unidirectional transformer is hard to model bi-directional sequences. Consequently, recent works have shifted focus towards paradigms like masked autoregressive (Li et al., 2024b; Chang et al., 2022; Yu et al., 2024c; Fan et al., 2024) and next-scale prediction (Tang et al., 2024; Tian et al., 2024; Han et al., 2024), which employ bidirectional attention in autoregressive models and demonstrate a superior choice for visual modeling. However, these approaches complicate visual generation and diverge from the traditional autoregressive paradigm, increasing the challenges for multi-modal unification. In contrast to existing works that modify the model to fit the data properties, we propose an alternative perspective: Can we instead instill a forward-dependency into the token sequence, to align with the powerful simplicity of decoder-only AR models?
@@ 2 Related work
Autoregressive visual generation. Inspired by GPT-style language models (Achiam et al., 2023; Touvron et al., 2023a; b), early visual autoregressive methods (Van Den Oord et al., 2017; Lee et al., 2022) predict image tokens along the raster-scan order. However, this strict unidirectional paradigm struggles to effectively model the bidirectional spatial dependencies inherent in visual sequence, thereby limiting its performance ceiling. To resolve this fundamental conflict, research has pivoted to re-engineering the models or generation paradigms for stronger bidirectional context awareness. One prominent approach is masked autoregressive modeling, including MaskGIT (Chang et al., 2022) and MAR (Li et al., 2024b), which leverages bidirectional attention for multi-round parallel generation. Another category explores hierarchical or next-scale prediction, such as VAR (Tian et al., 2024), which proceeds autoregressively across scales yet bidirectionally within each scale. Furthermore, other efforts enhance global perception by maximizing all-directional generation probabilities (RAR, Yu et al. (2024b)) or adopting random generation orders (RandAR, Pang et al. (2024)). While effective, these approaches compromise the simplicity of the classical autoregressive paradigm by introducing complex generation mechanisms or training objectives. Unlike these strategies that adapt models to data, this paper focuses on reshaping the data itself. We propose a novel tokenizer that instills a causal dependency structure into the token sequence during the encoding stage, thereby enabling a standard decoder-only autoregressive model to unleash its full potential.
Image tokenizer, a type of Variational Autoencoder (VAE) (Kingma et al., 2013), plays an important role in visual generation domain (Wan et al., 2025; Wu et al., 2024; Wei et al., 2024; Tang et al., 2025; Peng et al., 2025). By mapping high-dimensional image pixels to a low-dimensional latent space, the image tokenizer significantly enhances the training efficiency of generative models. Among them, VQ-VAE (Van Den Oord et al., 2017) utilizes quantization techniques to discretize continuous latent features for autoregressive image generation. Subsequently, VQ-GAN (Esser et al., 2021) substantially improves visual fidelity by incorporating an adversarial loss. MagViT-v2 (Yu et al., 2024a) and FSQ (Mentzer et al., 2023) focus on algorithmic improvements to efficiently scale the codebook size. Besides, TiTok (Yu et al., 2024c) proposes a transformer-based tokenizer, encoding 2D image patches into a 1D sequence. FlexTok (Bachmann et al., 2025) explores multi-granularity semantic encoding, ranging from coarse to fine-grained. Despite numerous efforts, there has been limited exploration of how to design a tokenizer that simultaneously achieves high-quality reconstruction while being more conducive to subsequent autoregressive generation. Bridging this critical gap is the core goal of our work.
@@ 3.2 AliTok tokenizer
The reconstruction loss for the first stage, follows standard practices (Yu et al., 2024c), combining a mean squared error (MSE) loss , perceptual loss , quantization loss (Zheng and Vedaldi, 2023) , and adversarial loss :
@@
Retraining a bidirectional decoder. In the second stage, we freeze the encoder and codebook of the tokenizer and retrain a bidirectional decoder to improve detail consistency. As shown in Fig. 4(d), in addition to converting the attention mechanism in the decoder, we integrate 64 buffer tokens (Li et al., 2024b) to enhance modeling capabilities by increasing computational load. Meanwhile, the loss previously imposed on prefix tokens is removed to allow the decoder to focus on reconstruction quality. This two-stage training strategy enables our AliTok tokenizer to not only produce generation-friendly encoded tokens, but also maintain good reconstruction performance.
@@ 3.3 Autoregressive model
Our decoder-only autoregressive architecture follows the standard design in LlamaGen (Sun et al., 2024), which uses RMSNorm (Zhang and Sennrich, 2019) for pre-normalization and applies 2D rotary positional embeddings (RoPE) (Su et al., 2024) at each layer. Building upon LlamaGen, we make only a few modifications. First, since our method has 16 extra prefix tokens and needs to model 272 tokens, 2D RoPE cannot be used directly. Therefore, we use 1D RoPE for the prefix tokens and 2D RoPE for the remaining 256 tokens. Secondly, we introduce the QK-Norm operation (Team, 2024) in the attention module to stabilize the large-scale model training, aligning with recent multi-modal autoregressive models (Xie et al., 2024; Ma et al., 2025; Team, 2024). We do not conduct extensive exploration of autoregressive models and only use the conventional architecture to verify the effectiveness of the proposed tokenizer.
@@ 4.1 Experimental Setup
Models and evaluations. The design of the proposed tokenizer is based on TA-TiTok (Kim et al., 2025), utilizing ViT-B (Yu et al., 2024c) for the encoder and ViT-L for the decoder. We set the vocabulary size to 4096 and adopt the online feature clustering method (Zheng and Vedaldi, 2023) to ensure that the codebook utilization is 100%. For simplicity, we call the autoregressive models trained using our AliTok tokenizer as AliTok-B/L/XL, comprising three different sizes, more details are provided in Appendix A.4. Following common practice, FID (including rFID and gFID) (Heusel et al., 2017), IS (Salimans et al., 2016), Precision, and Recall are adopted as metrics. We generate 50,000 images to test the gFID and adopt the evaluation code from RAR (Yu et al., 2024b). For rFID, we use the ImageNet-1K val set for evaluation, consistent with other methods (Sun et al., 2024; Yu et al., 2024c; Yao et al., 2025). During the test, KV-cache is employed to enhance sampling speed.
Dataset and training details. We train our image tokenizer from scratch on ImageNet-1K (Russakovsky et al., 2015). The tokenizer is trained for 600K steps in the first stage and 300K steps in the second stage on 32 A800-80G GPUs. For the autoregressive model training, we employ tencrop (Szegedy et al., 2015) for data augmentation and cache (Li et al., 2024b) the encoded results of the tokenizer to reduce training time. We train the base model and large model for 800 epochs and the XL model for 400 epochs, with a batch size of 2048 and a learning rate of 4e-4. For the first 100 epochs, the learning rate is linearly warmed up, then decayed using a cosine decay schedule down to 1e-5, following RAR (Yu et al., 2024b). During training, class conditioning is randomly dropped with a probability of 0.1 to support the use of classifier-free guidance (cfg) (Ho and Salimans, 2022).
@@ 4.2 Main Results
Sampling efficiency comparison. Beyond its excellent generation quality, AliTok also demonstrates a significant advantage in sampling efficiency benefiting from the use of KV-cache. As shown in Table 2, with similar gFID scores, AliTok-L improves throughput by 33.7 and 2.0 compared to MAR-H (Li et al., 2024b) and RAR-XXL (Yu et al., 2024b), respectively. Compared to the state-of-the-art diffusion model LightningDiT (Yao et al., 2025), our AliTok-XL requires less than 10% of the time needed by LightningDiT to produce an image, establishing a substantial improvement in image generation efficiency.
Generation visualization. We further show the generation results under different model sizes in Fig. 5. It can be observed that even the smallest model is capable of producing high-quality visual results and reasonably generating complex structures, such as the coral reef and altar. As the model size increases, the generated images exhibit enhanced detail and more intricate structures, illustrated by the realistic texture of the feathers of the parrot, the fine-grained structure of the space shuttle and the altar, demonstrating a superior level of realism in image quality. More results generated by the different model sizes are provided in the Appendix (Fig. 14, 15, 16, 17, and 18).
@@
Reconstruction performance. We report the reconstruction performance of various tokenizers (Yu et al., 2024c; Zheng and Vedaldi, 2023; Xiong et al., 2025) in Table 4. By introducing a powerful bidirectional decoder in the second stage, AliTok significantly improves the rFID to 0.86. This result remains competitive with GigaTok-B-L (rFID: 0.81), particularly given that the latter relies on a substantially larger vocabulary size (16384 4096). Experiments demonstrate that our proposed AliTok successfully achieves an ideal balance between the goals of generation-friendly and high-fidelity through its ingenious decoupling design.
@@ 4.3 Ablation Studies
| Method | #params | gFID | IS |
| REPA | 675M | 2.08 | 274.6 |
| MAR-L | 481M | 1.73 | 279.9 |
| AliTok-L | 318M | 1.39 | 295.3 |
Scalability in high-resolution generation. To verify the scalability of our method on the high-resolution image generation task, we trained AliTok on the ImageNet dataset at 512512 resolution. For this experiment, both the tokenizer and autoregressive model are fine-tuned from pre-trained weights at 256 resolution, where the embedding of tokenizer is linearly interpolated to match the shape. The generation results are shown in Table 5. Our AliTok-L achieves a gFID of 1.39 and an IS of 295.3, significantly surpassing MAR-L (Li et al., 2024b) and REPA (Yu et al., 2024d). The result highlights the strong scalability and robustness of our AliTok. Fig. 9 provides generated samples, visually confirming the high quality of the outputs.
| Size | rFID | gFID | IS | Loss | Acc |
| 4096 | 0.86 | 1.28 | 306.3 | 4.67 | 12.2% |
| 16384 | 0.81 | 1.30 | 304.9 | 6.42 | 4.3% |
Scalability in codebook size. To further validate the adaptability of the AliTok framework to larger discrete vocabularies, we conducted an experiment expanding the codebook size from 4096 to 16384. As shown in Table 6, the 16384 codebook achieves a lower reconstruction FID (from 0.86 to 0.81), indicating a higher upper bound for potential generation quality. On the generation side, after training our XL-sized model for 400 epochs, the generative performance with the 16384 codebook (gFID: 1.30) is slightly inferior to that of the 4096 codebook (gFID: 1.28). We attribute this temporary performance gap to the increased difficulty of the next-token prediction task within a 4 larger vocabulary space, which is clearly evidenced by the sharp drop in training accuracy and the rise in loss. Consequently, we posit that with extended training epochs or an increase in model parameters, the tokenizer with 16384 codebook size is expected to achieve better generation performance.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants 62225207, 62436008, 62306295, and 62576328. The AI-driven experiments, simulations and model training were performed on the robotic AI-Scientist platform of Chinese Academy of Sciences.
@@ A.1 Limitation and future work
Benefiting from the AliTok tokenizer, our 662M AR model reaches a significantly high training accuracy of 12.2%. In this context, the reconstruction performance of the discrete tokenizer becomes a critical bottleneck for the generative performance. To test the limiting value of gFID that can be achieved under this tokenizer, we randomly select 50 real images per class from the training set and calculate the gFID after reconstruction (differently from rFID evaluation), obtaining a FID of 1.15. This implies that even if the accuracy in the generation phase is close to 100% (which is not possible), the generation performance still cannot reach the upper limit of 1.15. This experimental result explains the phenomenon observed in Fig. 6, where the training loss and error rate of AliTok-XL continue to decrease steadily, yet the gFID metric struggles to achieve further improvement. While our experiments with a 16384-size codebook have demonstrated a higher ceiling for reconstruction quality (as shown in Table 5), they also revealed an increased learning challenge for the autoregressive model. This paper does not explore generative models larger than our XL variant, and we leave the comprehensive exploration, such as training larger-scale models or extending the training duration on the 16384 codebook to fully unlock its generative potential to future work.
@@ A.4 Experiments
| Tokenizer | TiTok-B | TiTok-L | GigaTok-B-L | AliTok (Stage 1) | AliTok (Stage 2) |
| Codebook size | 4096 | 4096 | 16384 | 4096 | 4096 |
| Codebook dimension | 12 | 12 | 8 | 32 | 32 |
| Training epochs | 200 | 200 | 200 | 120 | 60 |
| #Param | 204.8M | 641.1M | 621.6M | 389.8M | 390.0M |
| Total batch size | 256 | 256 | |||
| Base learning rate | 1e-4 | 1e-4 | |||
| Minimum learning rate | 1e-5 | 1e-5 | |||
Tokenizer training specifications. This subsection details the training configurations and model specifications for AliTok and other tokenizers, summarized in Table 8. As shown, our comparative experiments are conducted under fair conditions, with aligned batch sizes and learning rate schedules, ensuring the validity of our system-level performance analysis.
@@
| Tokenizer | Tokens | Size | PSNR | SSIM | LPIPS | rFID | #Param |
| TiTok-L-32 | 32 | 4096 | 16.15 | 0.5098 | 0.3232 | 2.21 | 641.1M |
| TiTok-B-64 | 64 | 4096 | 17.39 | 0.5491 | 0.2539 | 1.70 | 204.8M |
| AliTok (Stage 1) | 272 | 4096 | 20.66 | 0.6571 | 0.1659 | 0.91 | 389.8M |
| AliTok (Stage 2) | 272 | 4096 | 20.97 | 0.6610 | 0.1573 | 0.86 | 390.0M |
| GigaTok-B-L | 256 | 16384 | 21.24 | 0.6873 | 0.1284 | 0.81 | 621.6M |
| AliTok (Stage 1) | 272 | 16384 | 21.08 | 0.6805 | 0.1367 | 0.84 | 390.2M |
| AliTok (Stage 2) | 272 | 16384 | 21.30 | 0.6851 | 0.1303 | 0.81 | 390.4M |
System-level comparison of reconstruction performance. To comprehensively evaluate the performance of our proposed AliTok on the image reconstruction task, we detail the performance of AliTok against other tokenizers across several standard reconstruction metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Learned Perceptual Image Patch Similarity (LPIPS), and reconstruction FID (rFID).
The results in Table 10 highlight AliTok’s competitiveness. In the comparison with GigaTok-B-L using a 16384 codebook, our method achieves on-par reconstruction fidelity, matching its rFID of 0.81, while using approximately 37% fewer parameters (390.4M 621.6M). This performance also validates our two-stage design, as Stage 2 significantly enhances the reconstruction quality from Stage 1 to reach this state-of-the-art level. This demonstrates that AliTok can maintain great reconstruction performance while simultaneously creating a generation-friendly token sequence.
@@
Visualization of codebook. We performed a t-SNE analysis to visualize the codebooks learned by a standard ViT-VQ (conventional approach in our paper) and AliTok. We encoded 200 images from a single ImageNet class with each tokenizer, and collected the utilized codebook vectors for 2D projection. The resulting visualization in Fig. 12 shows a clear topological divergence. The baseline’s feature space (blue) is diffuse and continuous, characterized by ambiguous cluster boundaries and significant feature entanglement. This suggests that the unconstrained bidirectional encoder creates non-causal dependencies across tokens to achieve high reconstruction fidelity. AliTok (red), by contrast, learns a well-structured discrete manifold of compact, clearly separated semantic clusters. This demonstrates that AliTok’s causal decoder constraint forces the encoder to learn more deterministic and disentangled representations. This structured and predictable feature distribution presents a more learnable target for the autoregressive model, which partly explains the significant improvement in generation accuracy observed with AliTok.
| Setting | Sampling Time (batch=64) | Speedup |
| AliTok-XL (w/o KV-Cache) | 107.76s | 1.0 |
| AliTok-XL (w/ KV-Cache) | 10.07s | 10.7 |
| Tokenizer | 256256 (batch=64, w/ cfg) | 512512 (batch=64, w/ cfg) | ||||
| Tokens | Sampling time | Memory | Tokens | Sampling time | Memory | |
| Baseline-XL | 256 | 9.20s | 14.81G | 1024 | 95.17s | 46.98G |
| AliTok-XL | 272 | 10.07s | 15.48G | 1056 | 100.67s | 48.33G |
| (+6.3%) | (+9.5%) | (+4.5%) | (+3.1%) | (+5.8%) | (+2.9%) | |
Detailed Analysis of Sampling Efficiency and Scalability. we provide a comprehensive analysis of the sampling efficiency of AliTok, explaining the primary driver behind its significant speed advantage and detailing the scalability of our approach with respect to the prefix tokens. First, AliTok’s 10 speedup over masked models and diffusion-based models primarily stems from its full compatibility with Key-Value (KV) caching. In sequential generation, KV-caching stores past states to avoid re-computation at each step. Our ablation study in Table 11 confirms its dominant role: disabling KV-caching skyrockets inference time from 10.07s to 107.76s.
Regarding scalability, we further quantified the overhead of additional prefix tokens at higher resolutions. As shown in Table 12, the relative impact of these tokens decreases significantly as resolution increases. For instance, scaling from 256256 to 512512 reduces the extra sampling time overhead from 9.5% to 5.8%. This favorable trend occurs because prefix tokens grow linearly (O(W)) while total tokens grow quadratically (O(WH)), making the overhead asymptotically negligible.
Generalization to Multimodal Text-to-Image Generation. To investigate AliTok’s generalization capability, we conducted a preliminary text-to-image experiment. Following the unified paradigm of Janus-pro (Chen et al., 2025), we built a 0.5B parameter model where a single autoregressive (AR) architecture handles both text understanding and visual generation. This model is based on LLaVA-OV-0.5B (Li et al., 2024a) and incorporates AliTok as its visual tokenizer. The model was trained on the CC12M dataset (Changpinyo et al., 2021) at 512512 resolution for only 30,000 steps on 64 A100 GPUs.
Fig. 13 shows a qualitative comparison against the 1B-parameter Janus-pro model (trained for 460K steps). Despite the significant disparity in model scale and training duration, our model demonstrates the ability to generate coherent images corresponding to the text prompts. This result serves as an initial validation of AliTok’s effectiveness and efficiency in a challenging multimodal autoregressive setting, highlighting its potential for broader applications.